Home - Search - Browse - Alphabetic Index: 0- 1- 2- 3- 4- 5- 6- 7- 8- 9
A- B- C- D- E- F- G- H- I- J- K- L- M- N- O- P- Q- R- S- T- U- V- W- X- Y- Z
More Details for 2008-03-13
ISS On-Orbit Status 03/13/08

Crew sleep cycle today: sleep 8:00am -4:30pm; wake 4:30pm -8:00am tomorrow.

STS-123/Endeavour docked smoothly last night at 11:49pm EDT at the PMA-2 (Pressurized Mating Adapter-2) port, 24 minutes behind schedule (due to loss of target lock by the CW {Continuous Wave} laser of the Shuttle's TCS {Trajectory Control Sensor} during the manual rendezvous phase, requiring manual lock re-acquisition). The RPM (R-Bar Pitch Maneuver) started at 10:26pm and was successfully completed at 10:34pm, with Whitson and Malenchenko taking 200-300 close-up photographs of Endeavour's bottom heatshield. The station now hosts ten occupants again as Mission 1J/A is underway. (At the point of docking, Peggy Whitson rang the traditional ship's bell and announced 'Endeavour landed!' The combined crew is comprised of ISS CDR Whitson, FE-1 Yuri Malenchenko, FE-2 LĂ(c)opold Eyharts, STS CDR Dominic Gorie, PLT Gregory Johnson, MS1 Robert Behnken, MS2 Mike Foreman, MS3 Takao Doi (Japan), MS4 Rick Linnehan, and MS5/FE-2-16 Garrett Reisman who replaces Eyharts as FE-2, as the latter returns on the Endeavour as MS-5.)

After the docking, the station was reoriented as usual to minimize the risk of micrometeoroid/debris impacts upon the Shuttle (-x-axis in velocity vector, +z-axis in local vertical, i.e. Shuttle trailing).

After hatches opened at 1:36am, the new crew was welcomed aboard the ISS and given the mandatory 25-min. safety briefing.

CDR Whitson then configuring the transfer equipment which allows the Shuttle to supply oxygen (O2) to the ISS PBAs (Portable Breathing Apparatus) in support of pre-EVA mask prebreathe for denitrogenation.

Before the airduct between ISS and the Shuttle was installed for ventilation, FE-1 Malenchenko performed the standard Orbiter atmosphere sampling using the Russian AK-1M sampler.

Yuri also switched USOS/RS (US Segment/Russian Segment) comm systems to their mated-flight mode.

Garrett Reisman assisted MS-3 Doi on the Orbiter deck in releasing the SLP-D1 PRLA (Spacelab Pallet Payload Retention Latch Assembly), after which PLT Johnson and MS1 Behnken maneuvered the SRMS (Shuttle Remote Manipulator System) to grapple and transfer the SLP, which carries the SPDM (Special Purpose Dexterous Manipulator), to the POA (Payload ORU Attachment) on the Robotics MBS/MT (Mobile Base System/Mobile Transporter) where it was berthed. (Dextre is too large to be assembled in the Shuttle cargo bay since it would interfere with cargo bay closure in an emergency. It will therefore be assembled at the MBS on the main truss, at Workstation 6 (WS6).)

SPDM Power Glitch: After successful SLP PDGF (Power & Data Grapple Fixture) connection to the POA, with electrical connectivity confirmed, repeated attempts to activate the SPDM PSU (Power Switching Unit) for powering on the 'keep-alive' heaters were unsuccessful. Engineers are suspecting a software timing issue, i.e., too little time for the Lab RWS (Robotics Work Station) between applying power to the PSU and establishing data comm. A software patch is in work to modify the timer settings to increase the time span. It is expected that the patch will be available on orbit in 24-48 hrs. There are no thermal constraints with the SPDM on the POA as long as all blankets remain installed. There exists a thermal limit (TTL) of ~5 days for the OTCM1 (ORU Tool Changeout Mechanism 1), but no constraints for OTCM2. If the patch is unsuccessful, an alternate power-up path would be tried from the SSRMS (Space Station Remote Manipulator System) by latching it to the SLP PDGF. If still unsuccessful, thermal limitation can be extended if EVA-1 & EVA-2 are modified to leave the current thermal protection blankets in place.

Other crew activities after the SLP transfer included:

Transfer of IELKs: Garrett Reisman transferred his IELK (Individual Equipment Liner Kit) from the Shuttle to the Soyuz TMA-11/15S crew return vehicle where Yuri Malenchenko installed it for the new FE-2. Leo Eyharts' IELK was pulled out and temporarily stowed for return on a future flight.
Sokol leak test: Malenchenko and Reisman also performed the standard leak checkout on Garrett's Sokol space suit which he would wear as a Soyuz passenger, then set it up for 'drying out'.
Food Frequency Questionnaire: Leo Eyharts filled out the regular FFQ (Food Frequency Questionnaire), his fourth, on the MEC (Medical Equipment Computer) (by means of these FFQs, NASA/ESA astronauts keep a personalized log of their nutritional intake over time on special MEC software. Recorded are the amounts consumed during the past week of such food items as beverages, cereals, grains, eggs, breads, snacks, sweets, fruit, beans, soup, vegetables, dairy, fish, meat, chicken, sauces & spreads, and vitamins);
EVA-1 Preparations: For tonight's first spacewalk by Linnehan and Reisman, Peggy Whitson worked in the Airlock (A/L), installing the REBAs (Rechargeable EVA Battery Assemblies) in the EMUs (checkout of REBA-powered suit equipment and tools was waived to get back on the timeline after the delayed docking), while Reisman configured the A/L E/L (Equipment Lock) and Peggy and Leo later set up the DCS-760 camera for the EVA and initiated charging the DCS battery (at least 3 hrs).
In preparation for tonight's transfer of the 'Kibo' JLP (Japanese Experiment Logistics Module-Pressurized Section) from the cargo bay to the Node-2 zenith (upper) CBM (Common Berthing Mechanism), Eyharts powered up the CBCS (Centerline Berthing Camera System) at the zenith hatch for a video checkout of the setup from the Orbiter Shuttle and later deactivated it again.

Yuri Malenchenko worked with the ground on activating of the Elektron oxygen generator at 32 amps, first pressurizing the BZh Liquid Unit with N2 (nitrogen) via laptop and later monitoring the external temperature of its secondary purification unit (BD) for the first 10 minutes of operations to ensure that there was no overheating. (During nominal operations a gas analyzer is utilized to detect hydrogen (H2) in the O2 line (which could cause overheating) but is not included in the control algorithm until 10 minutes after Elektron startup. A new EMI filter, recently (2/9) installed on the Elektron's current stabilizer (FPP ST-64), is designed to prevents RFI (radio frequency interference) with the ATV/Automated Transfer Vehicle).)

Before sleep time, Malenchenko terminated the regeneration process on absorbent bed #1 of the Russian BMP (Harmful Impurities Removal System). Bake-out of bed #2 will follow tonight after wake-up (~5:03pm). (Regeneration of each of the two cartridges takes about 12 hours and is conducted only during crew awake periods.)

Working from the voluntary 'available time' suggestions list, Yuri Malenchenko conducted the regular daily checkup on the Japanese experiment GCF-JAXA (Granada Crystallization Facility) in the Russian TBU incubator, maintained at +20 degC, including a temperature check on its ART (automatic temperature recorder).

A second task list item for the FE-1 this morning was the daily monitoring, picture-taking and downloading for the BIO-5 Rasteniya-2 ("Plants-2") experiment which researches growth and development of plants (peas) under spaceflight conditions in the Lada-12 greenhouse from IBMP (Institute of Bio-Medical Problems {Russian: IMBP}).

For both crews, currently asleep, workday begins later today at 4:30pm EDT. Rick Linnehan and Garrett Reisman are on 'Campout' (nachalo desaturatsiy = desaturation start) in the A/L CL (Crewlock), after hatch closure this morning at ~7:40am. The two spacewalkers performed PBA (Portable Breathing Apparatus) mask prebreathe for denitrogenation, while readying their tools & equipment, then depressed the CL from 14.7 to 10.2 psi for their sleep period. (For the Campout, fresh METOX (Metal Oxide) canisters are installed in the A/L for CO2 control.)

The CL hatch will be cracked (i.e., temporarily repressurized) at ~5:10pm tonight for a hygiene break/with mask prebreathe for Reisman and Linnehan. Around 5:55pm, the hatch will be closed again for EVA preps in 10.2 psi, followed by EMU purge & prebreath. Afterwards, Peggy and Mike will support CL depressurization, and EV1 & EV2 will egress around 9:23pm. (EVA-1 will last approximately 6.5 hrs, ending at ~4:00am (ISS crew sleep begins at 8:00am). Its major objectives are: (1) prepare for JLP unberthing in cargo bay; (2) install OTCM2 with the Node-2 PDGF-based SSRMS, (3) remove and install OTCM1 (with SSRMS); (4) prepare for SPDM assembly, leaving thermal covers intact; (5) cleanup & ingress.)

ATV 'Jules Verne' Update (Flight Day 4): All ATV systems continue to perform nominally. Yesterday, ATV successfully performed two additional maneuvers to test PDE (Propulsion Drive Electronics) thruster configuration B (PDE-3 & -4). The first burn, TP1, began at 9:12:23am EDT, with a delta-V of 6.21 m/s; the second, TP2, at 10:01:29am, with a delta-V of 6.12 m/s. Both PDEs and He (helium) pressure regulation performed nominally, and ATV-CC/Toulouse reported that the boosts were nominal. Based on ATV GPS (Global Positioning System) data, specialists estimate that the post-maneuver state was within 100 m in semi-major axis of ATV-CC prediction. In addition, NASA-GSFC successfully performed coherent TDRSS tracking before and after the maneuver that confirmed the maneuver results. Per the nominal plan, at 1:55pm ATV performed an attitude slew (for a Helium purge), during which a thruster temperature differential alarm was annunciated that resulted in PDE-2 being switched off by the ATV FDIR (Fault Detection, Isolation & Recovery) system, although the ATV completed the slew maneuver nominally. Toulouse believes the cause to be FDIR thresholds set too low, and efforts are underway to update the limit settings, after which the PDE will be re-integrated into the propulsive system. At this time, there is no reason to believe that this anomaly is related to the earlier problem with the Helium pressurant system that resulted in PDE-2 being shut down.

Columbus Update: BLBGB (BIOLAB Glovebox) remains inoperable; one of its two rotors is jammed, rendering four ECs (Experiment Containers) on the other centrifuge inaccessible. The remaining four ECs (of 8 total) are recovered and secured in the TCU (Thermal Control Unit) for return. FSL (Fluid Science Laboratory) is not working, pending restoration with a new bus connector. SOLAR is working but degraded by a technical problem. EuTEF: six of 9 experiments are running; three are in troubleshooting process.

ISS Crew Sleep Shift Planning: To synchronize the ISS crew's timeline with STS-123/1J/A docking and subsequent docked activities, Peggy's, Yuri's and Leo's wake/sleep cycle is undergoing a number of shifts which started on 3/11. For the next six days, the wake/sleep shift schedule is as follows (all times EDT):

FD4 Wake: 4:30pm (3/13) - 8:00am (3/14)
Sleep: 8:00am - 4:30pm (3/14)

Wake: 4:30pm (3/14) - 7:00am (3/15)
Sleep: 7:00am - 3:30pm (3/15)

Wake: 3:30pm (3/15) - 7:00am (3/16)
Sleep: 7:00am - 3:30pm (3/16)

Wake: 3:30pm (3/16) - 6:00am (3/17)
Sleep: 6:00am - 2:30pm (3/17)

Wake: 2:30pm (3/17) - 6:00am (3/18)
Sleep: 6:00am - 2:30pm (3/18)

Wake: 2:30pm (3/18) - 5:00am (3/19)
Sleep: 5:00am - 1:30pm (3/19)

ISS Orbit (as of this morning, 8:44am EDT (= epoch)):
Mean altitude -- 340.6 km
Apogee height -- 341.5 km
Perigee height -- 339.8 km
Period -- 91.35 min.
Inclination (to Equator) -- 51.64 deg
Eccentricity -- 0.000131
Solar Beta Angle -- 16.0 deg (magnitude increasing)
Orbits per 24-hr. day -- 15.76
Mean altitude loss in the last 24 hours -- 173 m
Revolutions since FGB/Zarya launch (Nov. 98) -- 53343

Home - Search - Browse - Alphabetic Index: 0- 1- 2- 3- 4- 5- 6- 7- 8- 9
A- B- C- D- E- F- G- H- I- J- K- L- M- N- O- P- Q- R- S- T- U- V- W- X- Y- Z
© 1997-2017 Mark Wade - Contact
© / Conditions for Use